OpenStudy (anonymous):

Write the equation 3x-2y=16 in slope-intercept form.

OpenStudy (anonymous):

3X-2Y=16 -2Y=3X+16 Y=-3/2X + -8

OpenStudy (anonymous):

_2y=-3x+16

OpenStudy (anonymous):

YEAH SORRY

OpenStudy (anonymous):

-2Y = -3X + 16 Y = 3/2X - 8 <-------- SLOPE-INTERCEPT FORM

OpenStudy (anonymous):

Thats in slope intercept form. Y=Mx+b

OpenStudy (anonymous):

THANK YOU! :)